### v2

# Towards Interpretability Without Sacrifice: Faithful Dense Layer Decomposition with Mixture of Decoders

Abstract
Multilayer perceptrons (MLPs) are an integral part of large language models, yet their dense representations render them difficult to understand, edit, and steer. Recent methods learn interpretable approximations via neuron-level sparsity, yet fail to faithfully reconstruct the original mapping–significantly increasing model’s next-token cross-entropy loss. In this paper, we advocate for moving to layer-level sparsity to overcome the accuracy trade-off in sparse layer approximation. Under this paradigm, we introduce Mixture of Decoders (MxDs). MxDs generalize MLPs and Gated Linear Units, expanding pre-trained dense layers into tens of thousands of specialized sublayers. Through a flexible form of tensor factorization, each sparsely activating MxD sublayer implements a linear transformation with full-rank weights–preserving the original decoders’ expressive capacity even under heavy sparsity. Experimentally, we show that MxDs significantly outperform state-of-the-art methods (e.g., Transcoders) on the sparsity-accuracy frontier in language models with up to 3B parameters. Further evaluations on sparse probing and feature steering demonstrate that MxDs learn similarly specialized features of natural language–opening up a promising new avenue for designing interpretable yet faithful decompositions. Our code is included at: https://github.com/james-oldfield/MxD/.
1 Introduction
One strategy for addressing concerns about large language models’ (LLMs) [1, 2, 3] behavior is via a bottom-up approach to understanding and controlling the network internals–developing models of how and where human-interpretable features are represented in LLMs and how they affect the output [4, 5, 6]. Such a mechanistic understanding has proved helpful for a number of issues relating to safety and transparency, from controlling refusal of harmful requests [7] to detecting generation of unsafe code [6] and latent model knowledge [8].
However, developing models of LLMs’ internals faces challenges due to the dense nature of their representations [9, 10]. Indeed, many studies have found that individual neurons in MLP layers encode multiple distinct concepts. Rather than human-interpretable features being neatly aligned with individual neurons, they are often distributed across many [11, 12]. As a result, it is not straightforward to cleanly isolate specific concepts of interest in the models’ latent token representations.
Traditionally, imposing constraints on model form has offered a way to instill more predictable properties or structure. Indeed, there is a rich history of success with constraints in machine learning: from parts-based representations through non-negativity [13, 14], to structure through low-rankness or assumptions on geometry [15, 16]. With the particular issues posed by dense representations in LLMs, specialization through sparsity has re-emerged as a dominating strategy for learning more interpretable representations. With prior work showing that sparser models both aid human explanation [17] and achieve higher scores on LLM-based auto-interpretability metrics [18, 19], sparsity is often used as a proxy for interpretability [20, 21]. To this end, many recent works–such as sparse autoencoders [22, 23, 6]–take inspiration from traditional sparse dictionary learning methodologies [24, 25], re-writing pre-trained LLMs’ activations as sparse, non-negative linear combinations of atoms in a learned overcomplete basis. However, as argued in [26], such approaches do not learn the functional mechanisms of LLMs’ layers, and their inherent post-hoc nature demands additional parameters and computation on top of the base models.
Figure 1: Units of specialization for sparse layer variants: Neuron-level sparsity of existing sparse MLPs [27, 26] (center) vs layer-level sparsity (right), which the proposed Mixture of Decoders (MxD) layer enables at scale. For GPT2, the dimensions are: $O=768$, $H^{*}=O\cdot 4$, $H\approx N\approx O\cdot 32$.
One alternative approach is to directly replace layers with more interpretable equivalents [28], such as with wide MLPs with sparsity constraints. Transcoders [27, 29, 30, 26] (TCs) are a recent example of this, training new MLPs to mimic the functional behavior of MLPs with sparse hidden units, which have recently been shown to also learn more interpretable features [26]. Thus, instead of relying on external post-hoc analysis, sparse MLP layers offer a way to distill specialized features directly into the model’s forward pass itself.
Both of the above methods for learning specialized features fall into the same category of what one may call ‘neuron-level sparsity’. Dictionary learning methods restrict the number of non-zero elements used from a learned dictionary, whilst sparse MLPs [27] limit the number of active rows used from a learned ‘decoder’ matrix. At its core, whilst this constraint is useful for interpretability, it is too restrictive–often heavily trading off accuracy for sparsity, poorly reconstructing the original model components [31, 28]. We argue that preserving the base models’ performance is a crucial component of sparse MLP layer approximations for the following two key reasons:
  1. 1.
    Model faithfulness: sparse layers that poorly approximate the original layers risk missing critical intricacies of the base models’ behavior or latent features [32]. Conversely, an accurate reconstruction (yielding similar downstream next-token loss) is some evidence that the combination of newly learned subcomputations faithfully emulates the base model.
  2. 2.
    Practical adoption: sparse layers that closely preserve base models’ performance are capable of replacing the existing MLPs, directly integrating specialized computation into the native forward pass. Otherwise, downstream use of the sparse layers’ features must run on top of the base models’ computation. This introduces additional inference-time cost to every forward pass, and restricts any analysis to post-hoc settings.
In this paper, we advocate for moving from neuron-level to layer-level sparsity (as illustrated in Figure˜1) to address this. We propose the Mixture of Decoders (MxD) layer to overcome the sparsity-accuracy trade-off through scalable, resource-efficient conditional computation. Rather than individual vectors, MxDs learn interpretable sublayers as atomic units of specialization. This faithfully mirrors the functional form of dense layer we wish to approximate, and allows MxDs to readily generalize to modern MLP variants (i.e., the Gated Linear Unit [33]).
At a technical level, MxDs are constructed via a flexible tensor factorization [34] with the Hadamard product [35]. Through their parameter efficiency, MxDs scale the number of specialized layers far beyond what is feasible with classic sparse mixture of experts (MoEs) [36], and recover prior adapter-based MoEs [37, 38] as a special case. Crucially, we prove that the proposed tensor factorization in MxDs leads to each ‘expert’ sublayer implementing a linear transformation with full-rank weights–allowing faithful reconstruction even under heavy sparsity. Empirically, we demonstrate that MxDs significantly outperform alternative sparse MLP layers such as Transcoders [27] and Skip Transcoders [26] on the sparsity-accuracy frontier. In addition to their faithfulness, MxDs remain competitive with the SOTA on interpretability metrics. Our contributions can be summarized as follows:
  • We propose Mixture of Decoders, an instance of a flexible class of parameter-efficient MoE through Hadamard product-factorized weight tensors.
  • We prove that each specialized MxD expert’s weights inherit up to the same rank as the original MLP’s decoder, providing faithful approximation even in very sparse models.
  • Across $108$ sparse layers in $4$ LLMs (with up to 33B parameters) MxDs (i) pareto-dominate existing techniques on the sparsity-accuracy frontier yet (ii) remain competitive on $34$ sparse probing and steering tasks, validating the interpretability of the learned experts.
2 Methodology
We first recall the technical details of language models’ MLP layers and existing approaches to sparse approximations in Section˜2.1. We then introduce the proposed MxD in Section˜2.2, outlining the attractive rank properties it inherits in Section˜2.3 and factorized implementation in Section˜2.4. We conclude with extensions to modern MLP layers in Section˜2.5.
2.1 Preliminaries
Let $\mathbf{x}\in\mathbb{R}^{I}$ be the pre-MLP latent representation of a specific token at a given layer. Omitting bias terms throughout for brevity, the GPT2-style MLP layer produces the output vector $\mathbf{y}\in\mathbb{R}^{O}$ as:
$\mathrm{MLP}(\mathbf{x})={\mathbf{D}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}}^{\top}\mathbf{z}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\in\mathbb{R}^{O},\quad\text{with }\mathbf{z}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\coloneqq\phi\big({\mathbf{E}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}}^{\top}\mathbf{x}\big)\in\mathbb{R}^{H^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}},$ (1)
[…]
Sparse approximations
One approach to learning interpretable features in MLPs is to train new, wider MLPs with sparse hidden units to reconstruct the original layer’s outputs [27, 26, 30, 29], reminiscent of dictionary learning techniques [25]. In general, sparse MLPs share the model form:
$\text{SMLP}(\mathbf{x})=\mathbf{D}^{\top}\mathbf{z}=\sum_{h=1}^{H}z_{h}\mathbf{d}_{h}\in\mathbb{R}^{O},\quad\text{with }\mathbf{z}\coloneqq\mathcal{S}\big(\mathbf{E}^{\top}\mathbf{x}\big)\in\mathbb{R}^{H},$ (2)
where $\mathcal{S}(.)$ is a sparsity-inducing function (such as the top-$K$ [23] activation used in this paper). Here, the dimensionality of sparse MLPs’ learnable weights $\mathbf{E}\in\mathbb{R}^{I\times H},\,\mathbf{D}\in\mathbb{R}^{H\times O}$ are set as $H\gg H^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}$ such that the hidden layer is significantly larger than that of the original MLP. The original post-MLP output vectors are approximated as a $K$-sparse, non-negative linear combination of the rows $\mathbf{d}_{n}$ of a newly learned decoder matrix. Whilst this model form has been shown to learn interpretable, specialized features $z_{h}$ in language models [27, 26], their poor reconstruction is of questionable faithfulness and limits their use as a layer replacement in practice.
2.2 Mixture of Decoders
We now detail the proposed Mixture of Decoders (MxD) layer, which overcomes the sparsity-accuracy trade-off by treating sparsely activating linear layers as the atomic unit of specialization. We approximate the original MLP with a conditional combination of $N$ linear transformations:
$\mathrm{MxD}(\mathbf{x})=\sum_{n=1}^{N}a_{n}(\mathbf{W}_{n}^{\top}\mathbf{z})\in\mathbb{R}^{O},$ (3)
where $\mathbf{a}\coloneqq\mathcal{S}\big(\mathbf{G}^{\top}\mathbf{x}\big)\in\mathbb{R}^{N}$ are sparse ‘expert coefficients’ from learnable gating matrix $\mathbf{G}\in\mathbb{R}^{I\times N}$, and $\mathbf{z}\coloneqq\phi\big(\mathbf{E}^{\top}\mathbf{x}\big)\in\mathbb{R}^{H}$ is the dense output from an encoder. Here, $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ is a third-order tensor of parameters collating all $N$ experts’ decoder weights $\boldsymbol{\mathcal{W}}(n,:,:)=\mathbf{W}_{n}\in\mathbb{R}^{H\times O}$. In MxDs, we use a large $N$ to scale the feature specialization, and set $H:=H^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}$ to match the original MLP’s smaller hidden dimension.
With the gate routing each token to just its top-$K$ experts, each $\mathbf{W}_{n}\in\mathbb{R}^{H\times O}$ receives a gradient signal from only a specific set of semantically similar tokens. This implicit clustering naturally leads experts to specialize in feature-specific subcomputations, while collectively covering the layer’s full functionality. MxDs in Equation˜3 also directly inherit the MLP layers’ original functional form, avoiding the need to impose sparsity and non-negativity constraints on the hidden units $\mathbf{z}\in\mathbb{R}^{H}$. However, MxD decoders naively require a prohibitive $NHO$ parameters–preventing $N$ from scaling to tens of thousands of specialized components. To achieve parameter-efficiency whilst retaining layer capacity for faithful layer approximation, we parameterize MxDs’ third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ specifically to yield full-rank expert weights, defined elementwise as:
$\boldsymbol{\mathcal{W}}(n,h,:)=\mathbf{c}_{n}*\mathbf{d}_{h}\in\mathbb{R}^{O},\qquad\forall\,n\!\in\{1,\ldots,N\},\;h\!\in\{1,\ldots,H\},$ (4)
where * is the Hadamard product [34, 35], and $\mathbf{c}_{n},\,\mathbf{d}_{h}\in\mathbb{R}^{O}$ are the rows of learnable weights $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Intuitively, $\mathbf{D}$ implements a base transformation modulated by the $N$ specialized units in $\mathbf{C}$. Additional technical motivation for this parameterization with tensor methods can be found in Section˜A.3. This brings MxDs’ parameter count down significantly to $O\cdot(N+H)$ from $NHO$ in Equation˜3 with $N$ full decoders. One can then vary $N$ to parameter-match sparse MLP layers. We next detail how this design (i) retains expressivity in each unit for faithful layer approximation under sparsity in Section˜2.3 and (ii) yields a simple forward pass in Section˜2.4.
2.3 MxDs are rank-preserving
[…]
Lemma 1 (Decoder rank preservation).
We can materialize linear expert $n$’s weight matrix as $\boldsymbol{\mathcal{W}}(n,:,:)=\mathbf{W}_{n}=\mathbf{D}\,\mathrm{diag}\left({\mathbf{c}}_{n}\right)\in\mathbb{R}^{H\times O}$. Assuming $\mathrm{diag}\left({\mathbf{c}}_{n}\right)\in\mathbb{R}^{O\times O}$ is a diagonal matrix with no zeros along its diagonal (and thus invertible), we then have
$\mathrm{rank}(\mathbf{W}_{n})=\mathrm{rank}(\mathbf{D}\,\mathrm{diag}({\mathbf{c}}_{n}))=\mathrm{rank}(\mathbf{D}).$
The proof is found in Section˜A.1, which first derives the matrix-valued expression for each expert from Equation˜4 and then applies a standard rank equality. At a sparsity level of $K$, each MxD output vector is a weighted sum of $K$-many linear transformations (each with potentially full-rank weights) of the dense hidden units $\mathbf{z}$. As a result, MxDs retain layer capacity even under high sparsity. Sparse MLPs’ hidden units have only $K$ non-zero elements in contrast–each output in Equation˜2 is therefore confined to a $K$-dimensional subspace of $\mathbb{R}^{O}$, potentially limiting the capacity of sparse MLPs to faithfully approximate the original mapping in the small $K$ regime desirable for interpretability (mirroring speculations by [26]). Further, whilst alternative soft linear MoEs achieve scalability through low-rankness [40], Lemma˜1 states that no such rank constraints are present in MxDs. For approximating existing MLP layers where low-rank assumptions may not hold, MxDs are consequently a more suitable class of conditional layer.
2.4 Factorized forward pass
Figure 2: Mixture of Decoders extends the base MLP/GLU layers with a conditional ‘expert’ branch, modulating the MLP’s outputs.
MxDs compute a linear combination of $N$ linear transformations of the dense vector. With the proposed Hadamard-factorized weights, this yields a simple implementation.
Lemma 2 (Hadamard-factorized MoE forward pass).
Let $\mathbf{z}\in\mathbb{R}^{H}$ and $\mathbf{a}\in\mathbb{R}^{N}$ denote the MLP hidden units and expert coefficients respectively. Further, denote the decoder matrices as $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$ parameterizing $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$. MxD’s forward pass can be re-written as:
$\mathrm{MxD}(\mathbf{x})=\sum_{n=1}^{N}a_{n}\big(\mathbf{W}_{n}^{\top}\mathbf{z}\big)=\big(\mathbf{C}^{\top}\mathbf{a}\big)\,*\;\big(\mathbf{D}^{\top}\mathbf{z}\big).$ (5)
The proof is found in Section˜A.2. We include a notebook at https://github.com/james-oldfield/MxD/blob/main/form-equivalence.ipynb showing the equivalence in PyTorch. Further, please see Section˜A.5 for a discussion of how the Hadamard factorization relates to prior parameter-efficient MoEs with element-wise scaling [37], and Section˜B.6 for performance/computational cost comparisons.
Table 1: Model formulations of related work: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}}\in\mathbb{R}^{I},\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\in\mathbb{R}^{O}$ are the pre- and post-MLP representations respectively, $\mathbf{z}$ are the hidden units, and $\mathbf{a}$ is the vector of the ‘expert coefficients’ for MxD. Model-specific encoders/decoders $\mathbf{E},\,\mathbf{D}$ map between the hidden units and output.
MLPs SAEs Transcoders Skip Transcoders MxDs
[3] [22] [27] [26] (Ours)
[…]
Sparse component None $\mathbf{z}=\mathcal{S}\left(\mathbf{E}^{\top}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\right)\in\mathbb{R}^{H}$ $\mathbf{z}=\mathcal{S}\left(\mathbf{E}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}}\right)\in\mathbb{R}^{H}$ $\mathbf{z}=\mathcal{S}\left(\mathbf{E}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}}\right)\in\mathbb{R}^{H}$ $\mathbf{a}=\mathcal{S}\left(\mathbf{G}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}}\right)\in\mathbb{R}^{N}$
2.5 Extending MxDs to GLUs
In contrast to methods imposing neuron-level sparsity [22, 27, 26], MxDs do not make assumptions about the base layer’s encoder architecture or activation function. As a result, MxDs readily generalize to alternative architectures such as the Gated Linear Units (GLUs) [33] used in recent LLMs [1, 2]. Recall that GLUs’ hidden units are computed as $\mathbf{z}_{\text{GLU}}=\psi(\mathbf{E}_{\text{GLU}}^{\top}\mathbf{x})*\big(\mathbf{E}^{\top}\mathbf{x}\big)\in\mathbb{R}^{H},$ with additional GLU parameters $\mathbf{E}_{\text{GLU}}\in\mathbb{R}^{I\times H}$ and GLU activation function $\psi$ (e.g., Swish [1]). By substituting in the GLU hidden representations, MxDs straightforwardly extend the GLU model form too:
$\mathrm{MxD}_{\text{GLU}}(\mathbf{x})$ $=\sum_{n=1}^{N}a_{n}\,\mathbf{W}_{n}^{\top}\bigg(\underbrace{\psi(\mathbf{E}_{\text{GLU}}^{\top}\mathbf{x})*\big(\mathbf{E}^{\top}\mathbf{x}\big)}_{\text{GLU hidden units}}\bigg)=\big(\mathbf{C}^{\top}\mathbf{a}\big)\;*\;\mathbf{D}^{\top}\big(\psi(\mathbf{E}_{\text{GLU}}^{\top}\mathbf{x})*\big(\mathbf{E}^{\top}\mathbf{x}\big)\big)$
where $\mathbf{a}\coloneqq\mathcal{S}\big(\mathbf{G}^{\top}\mathbf{x}\big)\in\mathbb{R}^{N}$ are the expert units, and $\mathbf{W}_{n}=\mathbf{D}\,\mathrm{diag}(\mathbf{c}_{n})\in\mathbb{R}^{H\times O}$ as before. For a technical discussion of GLUs and their relationship to MxDs, we refer readers to Section˜A.4–through the theoretical results developed in this paper, we show that GLU encoders themselves can be viewed as a mixture of rank-$1$ linear experts (in contrast to the rank-preserving MxDs).
3 Experiments
The experimental section in the main paper is split into two parts. Section˜3.1 first demonstrates how MxDs perform significantly better on the accuracy-sparsity frontier as sparse MLP layer approximations on 4 LLMs. We then demonstrate in Section˜3.2 that MxD’s features retain the same levels of specialization through sparse probing and steering evaluations. Thorough ablation studies, experiments with matrix rank, and comparisons to low rank MoEs are presented in Appendix˜B.
3.1 Sparse approximations of MLPs in LLMs
In this section, we perform experiments approximating LLMs’ existing feed-forward layers with sparse MLPs, establishing that MxDs better navigate the sparsity-accuracy frontier, more faithfully approximating the base models’ MLPs than the SOTA baseline methods.
Implementation details
We train on $4$ base models: GPT2-124M [3], Pythia-410m, Pythia-1.4b [41], and Llama-3.2-3B [1] with up to 8080k experts/features. We train all sparse layers on a total of 480480M tokens of OpenWebText [42], with learning rate 1e41e-4 and a context length of $128$, initializing the output bias as the empirical mean of the training tokens, and $\mathbf{D}$ in MxDs as the zero-matrix (following [26]). We vary $N$ in MxD layers to parameter-match Transcoders in all experiments, with parameter counts and dimensions shown in Table˜2. For Llama3.2-3B, we use the Swish-GLU variant of MxD and GELU-MLP MxDs for the other three models, matching the architectures of their base encoders. Through ablation studies in Section˜B.8 we show that MxDs using the GELU/GLU variants are much more accurate layer approximators than the ReLU variants. Full experimental details are included in Appendix˜D. Whilst we do not have the computational resources to show similarly thorough experiments on even larger models, we expect MxDs to scale just as well as sparse MLPs to models with tens of billions of parameters or more.
Table 2: Sparse layer parameters/dimensions: $H$ denotes the size of the layers’ hidden units and $N$ is the expert count. MxDs perform almost as many linear transformations as the baselines have features.
GPT2-124M Pythia-410M Pythia-1.4B Llama-3.2-3B
Model Params $H$ NN Params $H$ NN Params $H$ NN Params $H$ NN
Transcoders [27] 37.737.7M 24,57624{,}576 67.167.1M 32,76832{,}768 268.5268.5M 65,53665{,}536 604604M 98,30498{,}304
Skip Transcoders [26] 38.438.4M 24,57624{,}576 68.268.2M 32,76832{,}768 272.7272.7M 65,53665{,}536 614614M 98,30498{,}304
MxDs 37.737.7M $3072$ 21,49021{,}490 67.167.1M $4096$ 28,65828{,}658 268.4268.4M $8192$ 57,33057{,}330 604604M $8202$ 86,01586{,}015
Objective function
[…]
3.1.1 Results: sparsity vs faithfulness
Figure 3: Model cross-entropy loss preserved when replacing MLPs with Transcoders [27], Skip Transcoders [26], and MxDs, as a function of the number of active units $K$ (hidden neurons/experts). We highlight that MxDs have consistently lower loss at all levels of sparsity.
We train an exhaustive set of $60$ sparse MLP approximations across $4$ diverse LLMs with up to 33B parameters. We show in Figure˜3 the resulting downstream base model cross-entropy loss when using the trained sparse layers in place of the original MLPs. As can be seen, not only do the proposed MxD layers outperform Transcoders [27] notably, but model performance is similarly preserved at all sparsity levels in MxD layers. Please also see Figure˜9 for results with normalized MSE, where MxDs’ reconstruction errors are up to an order of magnitude smaller. Full results on additional layers are included in Section˜B.3 for $48$ more trained sparse layers.
The recent ‘Skip Transcoders’ (STCs) [26], introduce an additional $IO$ parameters with a skip connection $\mathbf{S}\in\mathbb{R}^{I\times O}$ mapping the input directly to the output with $\mathbf{y}\approx\mathbf{D}^{\top}\mathbf{z}+\mathbf{S}^{\top}\mathbf{x}$. STC layers thus have considerably more parameters (e.g., STCs on llama3.2-3B have 1010M more parameters than MxDs). Despite the smaller parameter counts, we find MxDs consistently outperform STCs on the sparsity-accuracy frontier, attesting to the benefits of MxDs’ model form.
3.1.2 Faithfulness in output space
Next, we perform experiments comparing the faithfulness of sparse layers as their computation propagates to the model output space. We sample $16$ future tokens with the base model and then measure how similar the same generations are when the target MLP layer is replaced with the sparse layers. We use $512$ text snippets from OpenWebText, and take the first 4 words of each as the initial prompts, generating $16$ future tokens after each prompt. We plot in Figures˜4(a) and 4(b) the percentage of the samples’ continuations that are identical in the original LLM and hooked LLMs up to $n$ future tokens ahead. We note that this is a rather punishing task–any small deviations quickly compound as $n$ grows. Despite this, we see that the MxDs match the future token generations far better than the baselines, exhibiting more faithfulness in model output space (as well as in latent space).
Please see qualitative examples of the first $8$ prompts and the subsequent ‘diffs’ (using Python 3’s difflib) of the generated tokens in the Appendix in Figures˜7 and 8.
(a) Pythia-410m.
(b) GPT2-124m.
Figure 4: Proportion of $512$ generated samples that contain $n$ predicted future words identical to the original model’s output when replacing the base LLM’s MLP layer with the sparse layers.
3.2 Feature evaluations
The accurate reconstruction of MxD models in Section˜3.1 provides some evidence that MxDs are faithfully emulating the original MLP layers’ functional mapping. However, for interpretability, we care equally about the extent to which the learned features correspond to specialized, human-interpretable concepts. We confirm that MxD’s features compete with the baselines quantitatively in two ways: through probing for known concepts in Section˜3.2.1 and by steering the model using the learned features Section˜3.2.2. For all experiments in this section, we use the $K=32$ models.
Shared experts and specialization
Interestingly, we find MxDs naturally learn a ‘shared’ expert performing a common base transformation–the remaining K1K-1 active experts are thus free to dedicate their capacity to modelling features unique to individual tokens. This emergent shared/private processing complements recent trends to use shared experts by design in MoEs [43, 44, 45, 46, 47] with [43] arguing this facilitates greater specialization. Furthermore, one may view the skip connection in STCs [26] as performing an analogous role to the shared expert. With MxDs, however, all units have the same high capacity to accurately learn separate subcomputation regardless of the frequency or rarity of features.
We also observe that our trained MxDs exhibit very few ‘dead’ experts, as shown in Section˜C.1, with many experts contributing actively. Furthermore, initial ablations in Section˜C.2 show that one can train MxDs without shared experts if desired, at small performance cost. Please see qualitative results of activated tokens for particular experts in Appendix˜E.
Figure 5: Highest F1 score probing for ‘news category’ [48] on individual features/experts. As expected, the MxDs remain competitive with the Transcoder baselines, outperforming TopK-SAEs.
3.2.1 Sparse probing with individual features/experts
One challenge is that the sparse layers learn features in an unsupervised manner. As pointed out in [23], we therefore do not know which high-level features we ought to expect the model to learn (or even whether they exist in the OpenWebText training data). Nonetheless, we can reasonably expect a useful unsupervised model to learn at least a handful of commonly occurring concepts and linguistic themes. We accordingly focus our evaluation on the relative abilities of the sparse models to learn features well-predicting a variety of binary features used in the literature.
Concretely, to quantify the extent to which sparse layer features reliably fire in response to common high-level, interpretable concepts of natural language, we adopt the experimental settings of [49, 23, 19], training binary probes on the individual units of specialization (sparse hidden units $z_{n}$ for TCs/SAEs and expert units $a_{n}$ for MxDs–all pre-activation). For probing of sample-level concepts, we mean-pool activations across all non-padding tokens [19]. We train separate probes on $100$ features with the largest mean difference between positive and negative activations, as per [49].
We perform experiments on all $24$ binary probing tasks in the SAEBench suite [19]. Four of which are shown in Figure˜5, plotting the best F1 score (on a held-out set) for news topic classification in a 1-vs-all setting [48]. As can be seen, there exist individual MxD expert units that are predictive of various categories of news articles, competitive with the baselines. We refer readers to Section˜B.7 for additional experiments on $20$ more sample-level probing tasks, $10$ token-level probing tasks, and experimental details.
3.2.2 Feature steering
Figure 6: Mean score along dimensions of ‘textual coherence’ and ‘steerability’ of text generated by steering with the first 100 features of the sparse layers. Each sample is scored by 2 LLM judges.
Specific features might reliably fire in response to interpretable patterns of the input, yet not contribute to the generation process. Here, we aim to test this functional role of features by steering the LLMs. We note that these experiments do not aim to establish TCs/MxDs as competitive with the SOTA for controllable LLM generation. Rather, we aim to validate that the learned features contribute mechanistically to the LLM’s forward pass in a predictable way.
Mechanisms for steering
Let $\lambda\in\mathbb{R}$ be a hyperparameter controlling the desired ‘strength’ of the model edit. For TCs, we hook the forward pass at the relevant layer to increase the presence of target feature $n$ with $\hat{\mathbf{y}}=\mathbf{y}+\lambda\mathbf{d}_{n}$. In contrast, MxDs can be steered with $\hat{\mathbf{y}}=\mathbf{y}+\lambda\cdot(\mathbf{W}_{n}^{\top}\mathbf{z})$. Intuitively, increasing the weight of an expert’s contribution in the forward pass modulates the token representation in the direction of the learned specialization.
Results
We perform steering with the first $100$ neurons/experts individually, using $\lambda:=100$ for all experiments. We generate a collection of $10$ synthetic outputs for each neuron, each string consisting of $32$ generated tokens to the prompt ‘‘Let’s talk about ’’. We then ask two LLMs [^1] We use gemini-2.0-flash and llama-4-scout-17b-16e-instruct as two independent LLM judges. to rate the collection of text along two dimensions separately: (1) the extent to which a shared concept, theme, or linguistic pattern is present throughout the generated collection of text, and (2) the grammatical fluency of the text (please see Section˜D.1 for the full prompt). As can be seen from the mean scores over the $100$ neurons shown in Figure˜6, MxDs are competitive with the baselines, exhibiting a similar trade-off between textual coherence and presence of concept as we expect.
4 Related work
Sparse decompositions
Learning sparse [50, 25], non-negative [51] features of a data signal has found many applications in computer vision [15, 52, 53, 54] and natural language processing [55, 56, 57], motivated by the pursuit of interpretable, parts-based representations [13, 14]. In transformer-based language models [3], similar variants have been proposed for post-hoc analysis; sparse autoencoders (SAEs) are a popular method that rewrites latent features as non-negative combinations of atoms in a learned overcomplete dictionary, imposing either soft sparsity penalties [6, 22, 31] or thresholding activations directly [23, 58, 59]. Recent work aims to sparsify the existing layers of pretrained LLMs, learning new MLPs with sparse hidden units [29] for circuit analysis [27] or more interpretable yet faithful computation [26, 60]. Despite the surge of interest in SAEs, many works are emerging drawing attention to their limitations–underperforming baselines for probing [61], unlearning [62], and steering [63], in addition to other pathologies [64, 32, 65, 66].
Conditional computation
One natural alternative to static fully connected layers is conditional computation [67, 68]. Tracing back to the early work of [69, 70], single dense layers are replaced with specialized subunits–conditional on the input–as a form of layer-level sparsity. The Mixture of Experts (MoE) architecture [36, 71, 72] is a prominent example of conditional computation, breaking the link between parameter count and FLOPs. Consequently, MoEs have seen rapid adoption in SOTA models in recent years–scaling to very large parameter counts [73, 74, 75, 76, 77]. For parameter-efficient instruction tuning [37] introduces conditional (IA)3 adapters [38], modulating the MLP hidden dimension with the Hadamard product. Our proposed formulation with factorized weight tensors yields ‘MoVs’ [37] as a less scalable special case (see Section˜A.5). In contrast, MxDs model the decoder output space directly for reconstruction, and also provide significantly more specialized units than [37], making MxDs more suitable for our goal of interpretability.
Whilst the primary focus of MoEs has been on their impressive capabilities, the literature has observed that individual experts often specialize in particular semantic patterns of the input data, despite not being trained to do so [78, 79, 43, 80, 81]. For example, many works find that data that are in some sense similar are routed to the same experts–specializing to object shapes [82], texture [83], image category [84], or semantic patterns in natural language [36]. In the context of large language models, this emergent property of specialization in MoEs has been a primary focus of recent work: from encouraging monosemantic experts [85] or sparsity amongst experts’ weights [86] to efficiently scaling the expert count for fine-grained specialization [40]. In contrast to these works exploring pre-training, we explore an efficient design of MoE to replace existing LLMs’ dense layers.
Interpretability by design
Whilst MxDs follow the paradigm of specialization-via-sparsity, a number of promising interpretable mechanisms have been recently proposed that do not rely on sparsity constraints. For example, recent work explores bilinear layers [87], combinatorial structure [88], and/or tensor networks [89] as alternative ways of designing interpretable architectures. Notably, [87] show how the Hadamard product between two linear transformations of the same input vector allows direct interpretation (such as analyzing input-output feature interactions). Whilst MxDs yield a similar Hadamard product forward pass, it involves two different input vectors, and establishes a theoretical equivalence to mixture of experts and conditional computation when sparsity is present in one of the operands (i.e., the expert coefficients).
5 Conclusion
In this paper, we showed the benefits of decomposing dense layers’ computations as a mixture of interpretable sublayers. We proposed the Mixture of Decoders (MxD) layer to achieve this at scale, proving that MxDs’ linear experts preserve the matrix rank properties of the original decoders. Experimentally, we showed MxDs significantly outperform on the sparsity-accuracy frontier when trained to replace dense MLP layers. Quantitative results on sparse probing and feature steering demonstrated MxDs nonetheless learn specialized latent features similarly to existing interpretability techniques. Crucially, MxDs reexamine the dominating neuron-level sparsity paradigm of popular techniques, providing evidence that specialization doesn’t have to come with such a high cost to model performance. We believe MxDs (and specialization at the layer-level more generally) are an important step towards sparsity without sacrifice, and hope future work continues to build interpretable mechanisms that better preserve model capabilities. We are excited about future work exploring MxDs (and mixtures of linear transformations more generally) in alternative settings, such as for cross-layer features or transformations [90].
Limitations
Our experiments show MxDs outperform on the sparsity-accuracy frontier on 4 diverse LLMs. Whilst we fully anticipate this trend to scale just as well as with sparse MLPs in even larger models, our experiments only provide direct evidence for LLMs with up to 3B parameters, given our limited resources. Furthermore, whilst the TopK activation can greatly reduce the decoders’ FLOPs, the large encoders in sparse MLPs and the gating function in MxDs remain an additional inference-time cost. Future work could explore hierarchical structures [85, 36] and/or efficient retrieval [91] for further reductions in FLOPs. Secondly, MoEs are prone to issues of expert imbalance [71], or collapse [92]. Through random weight initialization alone, we find MxDs in our experiments learn diverse experts without collapsing to the base decoder. However, we expect standard MoE load-balancing [93] or diversity losses [94] to be useful for MxDs should one need more explicit ways of encouraging expert diversity, or when training new interpretable MxD architectures end-to-end. With regards to feature evaluations, our steering experiments rely on LLMs as judges. Whilst this is commonplace in recent popular steering benchmarks [63], there is mixed evidence emerging about the reliability of all base models [95, 96]. Our experiments attempt to circumvent issues by reporting scores across two capable SOTA models, but we caution inferring too much into the absolute values of reported scores (rather than the relative performance between sparse layers).
Acknowledgments
James Oldfield is grateful to Demian Till for reviewing the draft and providing valuable feedback and suggestions, and would also like to thank Markos Georgopoulos, Benjamin Hayum, and Wisconsin AI Safety Initiative’s Safety Scholars for insightful discussions throughout the project. We are also grateful to the open-source Zulip platform for facilitating research discussion. Sharon Li is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Shawn Im is also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2137424. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Support was also provided by the Graduate School and the Office of the Vice Chancellor for Research at the University of Wisconsin-Madison with funding from the Wisconsin Alumni Research Foundation. Mihalis Nicolaou is supported in part by the TensorICE project (EXCELLENCE/0524/0407), implemented under the social cohesion programme “THALIA 2021-2027”, co-funded by the European Union through the Research and Innovation Foundation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The quantitative results in Section˜3.1 support the claims that the MxDs better navigate the sparsity-accuracy frontier in models up to 33B parameters. Section˜3.2 quantitatively supports the claims that the features are similarly specialized.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are stated explicitly in Section˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The proofs of Lemmas˜1 and 2 are provided in the appendix in Section˜A.2, and referenced from the main paper. It is made explicit in Lemma˜1 under exactly which technical conditions the rank property holds. A small summary of the proof of Lemma˜1 is given in the main paper, too.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, all details, configurations, and resources used to train the models are included in Appendix˜D. Implementations of the models are provided in notebooks explicitly linked to from the main paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: A link to the code to train the models is included in the submission.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details about the main training runs are provided in Table˜10, and the details about the sparse probing experiments are provided in Section˜B.7.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The largest model runs take on the order of multiple days per value of KK–thus, training multiple models is not feasible on our resource budget.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, this is provided in Table˜10, including the specific GPU types and wall-clock runtime.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: No violations of the code of ethics are made. In particular, the paper focuses on designing more transparent and interpretable mechanisms, and thus we don’t consider the paper to cause any direct negative societal issues. However, we acknowledge that there is often an inherent dual-use nature to much interpretability work, including our own. Furthermore, whilst our largest training runs take multiple days, each uses a single GPU, therefore the environmental impact is limited.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss how our design of increasingly performant yet interpretable models may provide better understanding, debugging, and editing of pre-trained models. We do not explicitly discuss the downsides of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The proposed work only explores existing pre-trained models that are openly accessible.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We include links to all used models’ huggingface pages in Table˜10.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets are released.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects are used.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects are used.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We use 2 LLMs via public APIs for our steering evaluation. The full prompt is provided in Section˜D.1, and the exact model endpoint names are provided for reproducibility.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Appendix A Proofs and additional technical results
A.1 Proof of rank equality
Proof of Lemma˜1.
We first derive the expression for expert $n$’s weight matrix $\mathbf{W}_{n}=\mathbf{D}\,\text{diag}(\mathbf{c}_{n})\in\mathbb{R}^{H\times O}$ and then show the rank equality that follows. First, recall that we have the third-order weight tensor defined as
$\boldsymbol{\mathcal{W}}(n,h,:)=\mathbf{c}_{n}*\mathbf{d}_{h}\in\mathbb{R}^{O},$
for matrices $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. We can express each element of the tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ in terms of elements of the two matrices as
$\boldsymbol{\mathcal{W}}(n,h,o)=c_{no}\cdot d_{ho}=(\mathbf{D})_{ho}\cdot c_{no}.$ (6)
Equation˜6 shows that for a fixed expert $n$, the $n^{\text{th}}$ row $\mathbf{c}_{n}\in\mathbb{R}^{O}$ essentially scales the columns of matrix $\mathbf{D}\in\mathbb{R}^{H\times O}$. This is equivalent to right-multiplying matrix $\mathbf{D}$ by a diagonal matrix formed from $\mathbf{c}_{n}\in\mathbb{R}^{O}$. Indeed, the $(h,o)$ entry of such matrix product is
$\left[\mathbf{D}\,\text{diag}(\mathbf{c}_{n})\right]_{ho}$ $=\sum_{i=1}^{O}(\mathbf{D})_{hi}\,\text{diag}(\mathbf{c}_{n})_{io}$ (7)
$=(\mathbf{D})_{ho}\,\text{diag}(\mathbf{c}_{n})_{oo}$ (8)
$=d_{ho}\cdot c_{no},$ (9)
since all off-diagonal terms (i.e., $i\neq o$) in Equation˜7 vanish and $\text{diag}(\mathbf{c}_{n})_{oo}=c_{no}$ by construction. Comparing Equation˜6 and Equation˜9 shows that, for every $h\in\{1,2,\ldots,H\}$ and $o\in\{1,2,\ldots,O\}$ we have
$\boldsymbol{\mathcal{W}}(n,h,o)=\left[\mathbf{D}\,\text{diag}(\mathbf{c}_{n})\right]_{ho}.$
Hence, indexing into the first mode of the tensor alone gives us the matrix-valued expression for expert $n$ as claimed:
$\boldsymbol{\mathcal{W}}(n,:,:)=\mathbf{W}_{n}=\mathbf{D}\,\text{diag}(\mathbf{c}_{n})\in\mathbb{R}^{H\times O}.$
Finally, a standard result in linear algebra [97] has that $\text{rank}(\mathbf{AB})=\text{rank}(\mathbf{A})$ for any $\mathbf{A}\in\mathbb{R}^{H\times O}$ and invertible matrix $\mathbf{B}\in\mathbb{R}^{O\times O}$. Since matrix $\mathrm{diag}\left({\mathbf{c}}_{n}\right)\in\mathbb{R}^{O\times O}$ is invertible by assumption in Lemma˜1, setting $\mathbf{A}=\mathbf{D}$ and $\mathbf{B}=\text{diag}({\mathbf{c}}_{n})$ yields the rank equality. ∎
A.2 Proof of MxD forward pass equivalence
Recall we have input vector $\mathbf{z}\in\mathbb{R}^{H}$, expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$, and layer weights $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$. The weights are defined in Equation˜4 element-wise through the Hadamard product * as
$\boldsymbol{\mathcal{W}}(n,h,:)=\mathbf{c}_{n}*\mathbf{d}_{h}\in\mathbb{R}^{O},\quad\forall\,n\!\in\{1,\ldots,N\},\;h\!\in\{1,\ldots,H\},$
for learnable parameters $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Lemma˜2 states that MxD’s forward pass can be equivalently expressed as
$\sum_{n=1}^{N}a_{n}\big(\mathbf{W}_{n}^{\top}\mathbf{z}\big)=\big(\mathbf{C}^{\top}\mathbf{a}\big)\,*\;\big(\mathbf{D}^{\top}\mathbf{z}\big).$
Proof of Lemma˜2.
The LHS can first be re-written as an explicit sum over the hidden dimension
$\hat{\mathbf{y}}=\sum_{n=1}^{N}a_{n}\big(\mathbf{W}_{n}^{\top}\mathbf{z}\big)=\sum_{n=1}^{N}\sum_{h=1}^{H}a_{n}\big(\mathbf{w}_{nh:}z_{h}\big)\in\mathbb{R}^{O}.$ (10)
Plugging in the definition of $\mathbf{w}_{nh:}\in\mathbb{R}^{O}$ from Equation˜4 then yields
$\hat{\mathbf{y}}$ $=\sum_{n=1}^{N}\sum_{h=1}^{H}a_{n}\big(\mathbf{w}_{nh:}z_{h}\big)$ (11)
$=\sum_{n=1}^{N}\sum_{h=1}^{H}a_{n}\big(\left(\mathbf{c}_{n}*\mathbf{d}_{h}\right)z_{h}\big)$ (12)
$=\bigg(\sum_{n=1}^{N}a_{n}\mathbf{c}_{n}\bigg)*\bigg(\sum_{h=1}^{H}z_{h}\mathbf{d}_{h}\bigg)$ (13)
$=\big(\mathbf{C}^{\top}\mathbf{a}\big)*\big(\mathbf{D}^{\top}\mathbf{z}\big),$ (14)
which is exactly the RHS of Equation˜5, showing the MxD forward pass is equivalent to the Hadamard product of $\mathbf{C}^{\top}\mathbf{a}$ and $\mathbf{D}^{\top}\mathbf{z}$. ∎
A.3 Intuition for weight parameterization through the lens of tensor methods
A second complementary way of viewing the MxD layer’s parameterization (and its full-rank properties) is through the lens of tensor methods [34]. A tensor-based motivation for MxD’s weight tensor parameterization and forward pass is presented in Section˜A.3.1 and Section˜A.3.2, respectively.
Notation and definitions
A brief primer is first included below, based on [34] (and can be safely skipped for those already familiar):
  • The mode-$n$ fibers of an $N^{\text{th}}$ order tensor $\boldsymbol{\mathcal{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ are the $I_{n}$-dimensional column vectors obtained by fixing every index except that of the $n^{\text{th}}$ mode (e.g., $\mathbf{x}_{:{i_{2}}{i_{3}}}\in\mathbb{R}^{I_{1}}$ are the mode-$1$ fibers of a third-order tensor $\boldsymbol{\mathcal{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$). Stacking all mode-$n$ fibers column-wise yields the so-called mode-$n$ unfolding $\mathbf{X}_{(n)}\in\mathbb{R}^{I_{n}\times\bar{I}_{n}}$, with number of columns given by the product of remaining dimensions $\bar{I}_{n}=\!\!\prod_{\begin{subarray}{c}t=1\\ t\neq n\end{subarray}}^{N}\!I_{t}$.
  • The Khatri-Rao product (denoted by $\odot$) between two matrices $\mathbf{A}\in\mathbb{R}^{I\times K}$ and $\mathbf{B}\in\mathbb{R}^{J\times K}$, is the column-wise Kronecker product (denoted by $\otimes$):
    $\mathbf{A}\odot\mathbf{B}:=\bigl[\,\mathbf{a}_{:1}\otimes\mathbf{b}_{:1}\;\cdots\;\mathbf{a}_{:K}\otimes\mathbf{b}_{:K}\,\bigr]\;\in\;\mathbb{R}^{(I\cdot J)\times K}$.
  • The mode-$n$ (vector) product of a tensor $\boldsymbol{\mathcal{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ with a vector $\mathbf{u}\in\mathbb{R}^{I_{n}}$ is denoted $\boldsymbol{\mathcal{X}}\times_{n}\mathbf{u}$ and has entries $(\boldsymbol{\mathcal{X}}\times_{n}\mathbf{u})_{i_{1}\ldots i_{n-1}i_{n+1}\ldots i_{N}}=\sum_{i_{n}=1}^{I_{n}}x_{i_{1}i_{2}\ldots i_{N}}\,u_{i_{n}}$.
A.3.1 MxD weight tensors through the Khatri-Rao product
MxDs construct the collective weight tensor through the Khatri-Rao product $\odot$ [34] of the two factor matrices $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Concretely, the mode-$3$ unfolding222which is simply a reshaping of a higher-order tensor into a matrix, arranging all $N$ expert matrices’ column vectors along the columns of a new matrix. of the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ in MxDs from Equation˜4 is alternatively given by:
$\mathbf{W}_{(3)}:=\left(\mathbf{C}\odot\mathbf{D}\right)^{\top}\in\mathbb{R}^{O\times(N\cdot H)}.$ (15)
Given that the factor matrices are learned end-to-end without constraints, they are likely of full column-rank, i.e. $\text{rank}(\mathbf{D})=\text{rank}(\mathbf{C})=O$ (as $N>O,\,H=4\cdot O>O$ in practice given the MLP layers’ larger bottleneck). Consequently, their Khatri-Rao product parameterizing the collective $N$ experts’ weights will be of maximum rank $O$ too, through Lemma 1 of [98]. As a result, parameterized this way, the $O$-dimensional fibers likely span the full output space.
A.3.2 Tensorized MxD forward pass
Furthermore, the layer’s forward pass can then be viewed as performing two tensor contractions between the third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ (collecting all $N$ experts’ $H\times O$-dimensional matrices) and expert coefficients $\mathbf{a}\in\mathbb{R}^{N}$ and hidden activations $\mathbf{z}\in\mathbb{R}^{H}$. This can be expressed in terms of the so-called mode-$n$ product (denoted by $\times_{n}$) [34] as follows:
$\hat{\mathbf{y}}$ $=\sum_{n=1}^{N}a_{n}\cdot\left(\mathbf{W}_{n}^{\top}\mathbf{z}\right)$
$=\sum_{n=1}^{N}a_{n}\sum_{h=1}^{H}\mathbf{w}_{nh}z_{h}=\sum_{n=1}^{N}\sum_{h=1}^{H}a_{n}z_{h}\mathbf{w}_{nh}$
$=\boldsymbol{\mathcal{W}}\times_{1}\mathbf{a}\times_{2}\mathbf{z}\in\mathbb{R}^{O}.$ (16)
A.4 GLU encoders are a mixture of rank-1 linear experts
Both the proposed MxDs and Gated Linear Units (GLUs) [33] share a similar functional form, using the element-wise product. However, there are crucially important differences between GLUs and MxDs that make both their interpretation and model capacity different.
In short, the technical results here in our paper show that GLUs’ encoder can be viewed as a linear mixture of expert layer with rank-1 experts. Furthermore, GLUs can be modified and extended to MxDs with two additions to their model form as detailed at the end of this subsection. First, recall that the GLU encoder [33] computes:
$\mathbf{y}_{\text{GLU}}=\psi(\mathbf{E}_{\text{GLU}}^{\top}\mathbf{x})*\big(\mathbf{E}^{\top}\mathbf{x}\big)\in\mathbb{R}^{H},$ (17)
for input vector $\mathbf{x}\in\mathbb{R}^{I}$, learnable weights $\mathbf{E}_{\text{GLU}},\,\mathbf{E}\in\mathbb{R}^{I\times H}$, and activation function $\psi(.)$. To transform Equation˜17 into the same model form as MxDs, we first pre-multiply the LHS by the identity matrix to match the MxD model form of Equation˜5, yielding:
$\mathbf{y}_{\text{GLU}}=\big(\mathbb{I}^{\top}\mathbf{a}\big)*\big(\mathbf{E}^{\top}\mathbf{x}\big),$ (18)
where $\mathbf{a}=\psi(\mathbf{E}_{\text{GLU}}^{\top}\mathbf{x})\in\mathbb{R}^{H}$ and $\mathbb{I}\in\mathbb{R}^{H\times H}$ is the $H$-dimensional identity matrix. Next, we can write this explicitly in terms of a linear MoE with expert weights $\mathbf{W}_{n}\in\mathbb{R}^{I\times H}$ as follows:
$\mathbf{y}_{\text{GLU}}$ $=\big(\mathbb{I}^{\top}\mathbf{a}\big)*\big(\mathbf{E}^{\top}\mathbf{x}\big)$ (19)
$=\sum_{n=1}^{H}a_{n}\big(\mathbf{W}_{n}^{\top}\mathbf{x}\big)$ (20)
$=\sum_{n=1}^{H}a_{n}\left(\mathbf{E}\,\mathrm{diag}\left((\mathbb{I})_{n}\right)\right)^{\top}\mathbf{x}\big),$ (21)
where $(\mathbb{I})_{n}\in\mathbb{R}^{H}$ is the $n^{\text{th}}$ row of the $H$-dimensional identity matrix (i.e. a one-hot vector with its only non-zero element at index $n$). We draw particular attention to how the $n^{\text{th}}$ expert’s matrix $\mathbf{W}_{n}=\mathbf{E}\,\mathrm{diag}\left((\mathbb{I})_{n}\right)\in\mathbb{R}^{I\times H}$ essentially picks out the $n^{\text{th}}$ column of $\mathbf{E}$, leaving all remaining H1H-1 columns as zero vectors. Therefore, GLU encoders compute a MoE with linear expert weights of (at most) rank 1. This relationship between GLUs and conditional computation is consistent with prior work interpreting individual GLU column vectors as experts [99]. Whilst GLUs’ encoders’ model form does not put any inherent restrictions on the total number of rank-1 terms that can contribute to the output, the sparsity necessary for specialization does.
We conclude this section by summarizing the two technical changes needed to transform GLUs into full-rank linear MoEs based on the Hadamard product:
  1. 1.
    Replace $\mathbb{I}$ in Equation˜18 with learnable, non-diagonal weight matrices for full-rankness.
  2. 2.
    Choose $\psi(.)$ to produce non-negative, sparse coefficients to encourage specialization through sparsity among the experts (for example, a softmax function, or a ReLU activation followed by TopK).
The first of the steps above provides full-rankness, whilst the second brings the sparsity and non-negativity needed for specialization. We include a notebook showing this connection in PyTorch at: https://github.com/james-oldfield/MxD/blob/main/glus-to-moes.ipynb.
A.5 Hadamard-factorized tensors generalize MoVs
Prior work [37] proposes to linearly combine $N$ many (IA)3 adapters [38] for parameter-efficient MoEs for instruction fine-tuning. The implementation results in a very similar functional form to the factorized forward-pass in MxDs. Interestingly, the Hadamard product parameterization of the third-order weight tensor in Equation˜4 provides a more general framework through which one can also derive MoVs’ model form, shedding light on the relationship to the proposed MxDs and their benefits. Concretely, factorizing the weight tensor instead along the second mode as $\boldsymbol{\mathcal{W}}(n,:,o)=\mathbf{c}_{n}*\mathbf{d}_{o}\in\mathbb{R}^{H}$ in our framework immediately recovers MoV [37] as a special case. In particular, in contrast to the MxD in Section˜A.3 whose weight tensor can be parametrized equivalently through its mode-$3$ unfolding [34], MoV’s implicit weight tensor can be given in terms of its mode-$2$ unfolding in terms of a similar Khatri-Rao product of two factor matrices.
Instead, MoVs in analogy would yield expert weights by pre-multiplying $\mathbf{D}$ as: $\mathbf{W}_{n}=\mathrm{diag}(\mathbf{c}_{n})\,\mathbf{D}\in\mathbb{R}^{H\times O}$ for much larger $\mathbf{C}\in\mathbb{R}^{N\times H}$). Due to $H\gg O$, our proposed MxD formulation yields around $4\times$ the number of specialized units as MoVs with the same parameter budget (yet MoVs’ experts are of no higher rank than MxDs’), making MxDs a much more suitable and efficient class of layer for our goal of scalable specialization. We therefore see that the proposed lens of tensor methods for unification provides valuable insights about how to design more interpretable layers with the minimum trade-off to capabilities.
A.6 TopK activation function
Formally, for activations $\mathbf{z}\in\mathbb{R}^{H}$ (possibly with the ReLU activation applied), the TopK activation function can be formulated elementwise as the following:
$\text{TopK}(\mathbf{z})_{h}=\begin{cases}z_{h},&\text{if }z_{h}\geq\tau_{K}(\mathbf{z})\\ 0,&\text{otherwise}\end{cases},$
where $\tau_{K}(\mathbf{z})\in\mathbb{R}$ returns the value of the $K^{\text{th}}$ largest element of vector $\mathbf{z}\in\mathbb{R}^{H}$. Intuitively, the TopK activation retains only the largest $K\leq H$ elements of the $H$ neurons, setting the rest to zero. We note here that the definition above does not handle ties (although this is very unlikely in practice, given the activations are continuous).
Appendix B Additional quantitative results and ablations
B.1 Faithfulness in output space
Experiments in Section˜3.1.2 of the main paper report the number of future predicted words that are identical from the original model and from the LLM with the sparse layer replacements. We also show qualitative examples of the first $8$ prompts and the subsequent ‘diffs’ (using Python 3’s difflib) of the generated tokens in Figures˜7 and 8–we see MxDs’ superior ability to preserve model functionality as it propagates through to the output space of future tokens.
Figure 7: Pythia-410m: The first few generated tokens from the base model (‘GT’) and the corresponding tokens from the model when the sparse layers replace the target MLP. Red denotes tokens that are removed, orange denotes newly inserted tokens, and green denotes matching tokens.
Figure 8: GPT2-124m: The first few generated tokens from the base model (‘GT’) and the corresponding tokens from the model when the sparse layers replace the target MLP. Red denotes tokens that are removed, orange denotes newly inserted tokens, and green denotes matching tokens.
B.2 Additional reconstruction metrics
To highlight the scale of difference in the reconstructions between MxDs and the baselines, we also plot in Figure˜9 the normalized MSE at the end of training for all models and LLMs. At the smallest values of $K$ (which we care about most for interpretability), MxDs’ normalized MSE is up to an order of magnitude smaller than Transcoders’.
Figure 9: Normalized MSE at the end of training Sparse MLP layers, as a function of the number of active units (i.e., hidden neurons vs experts); with differences as large as an order of magnitude in error.
B.3 Results on additional layers
We also fully train all models and baselines (with 4 different values of $K$) on different target layers for each model. The results are shown in Figure˜10 for $48$ additional trained layers for the same setup in the original paper, using different colours to highlight that these are new results. As can be seen, the same trend holds: MxDs significantly outperform the baselines at small $K$ in all LLMs.
Figure 10: Additional layer results: model cross entropy loss preserved when replacing MLPs with Transcoders [27], Skip Transcoders [26], and MxDs, as a function of the number of active units (hidden neurons/experts). These results complement those in the main paper, but here we train a new set of additional models on different layers.
B.4 Initial results on Gemma2-27B
In the main paper, we perform experiments on base models with up to 3B parameters. For initial evidence of scalability to even larger models, we perform here additional large-scale experiments on the 27B Gemma2 model (in half-precision, and with a smaller batch size to fit into memory). We see MxDs continue to outperform on the sparsity-accuracy frontier, as shown by the (normalized) reconstruction loss in Table˜3.
Table 3: Normalized MSE ($\downarrow$) on Gemma2-27B, Layer 20, $K=32$, partially trained for 50k iterations. To fit in memory, the model is loaded in half precision. We use $1/8$ the batch size and # buffers stored, and $1/2$ the multiplier on the number of latent features (with values of 4, 16, and 16, respectively).
Iterations 10k 20k 30k 40k 50k
Transcoder
0.147 0.132 0.128 0.123 0.119
Skip Transcoder
0.128 0.107 0.102 0.100 0.093
MxD
0.098 0.096 0.093 0.086 0.069
B.5 Expert rank
This section concerns the matrix rank of the linear experts in parameter-efficient MoEs. We first compare to low-rank MoEs in Section˜B.5.1 to demonstrate the benefits of full-rankness, and then follow up in Section˜B.5.2 by confirming that the learned MxD expert ranks are close to maximum in the trained models.
B.5.1 Comparisons to low-rank MoEs
[…]
$\mu\text{MoE}(\mathbf{x})$ $=\sum_{n=1}^{N}\sum_{h=1}^{H}\sum_{r=1}^{R}a_{n}z_{h}\,\mathbf{D}(r,h)\cdot\mathbf{C}(r,n)\cdot\mathbf{W}(:,r)\in\mathbb{R}^{O},$ (22)
where $\mathbf{C}\in\mathbb{R}^{R\times N},\,\mathbf{D}\in\mathbb{R}^{R\times H},\,\mathbf{W}\in\mathbb{R}^{O\times R}$ are the learnable low-rank terms of the implicit third-order tensor parameterizing all $N$ collective experts’ weights.
We match the MxD experimental configuration as closely as possible for a fair comparison. For the encoders, we mirror MxDs and use the GELU activation function, which we find through ablations in Section˜B.8 to perform the best. We initialize the parameters the same as MxDs and Skip Transcoders: we use the standard PyTorch linear layer initialization for $\mathbf{D},\,\mathbf{C}$ (and the encoder layers), and initialize $\mathbf{W}$ as the zero matrix.
[…]
As a motivating example, for why SparseMoEs and SoftMoEs are not practical: SparseMoEs [36] and SoftMoEs [70] require $2.16$ trillion parameters for a single layer, for the same 8686k experts we use for Llama-3.2-3B. This is orders of magnitude more parameters than the entire base network itself, making it prohibitively costly for SparseMoEs to scale to sufficiently high expert counts.
B.5.2 MxD empirical expert rank
Next, we show experimentally that the learned experts’ matrices $\mathbf{W}_{n}=\mathbf{D}\,\text{diag}({\mathbf{c}}_{n})\in\mathbb{R}^{H\times O}$ are very nearly full-rank in practice, corroborating the properties of expert matrices shown theoretically in Lemma˜1. We compute the mean ‘normalized rank’, which we take for MxDs to be the empirical matrix rank of the learned expert’s weights, over the maximum possible rank given the dimensions:
$\frac{1}{N}\sum_{n=1}^{N}\frac{\text{rank}\left(\mathbf{W}_{n}\right)}{\min\{H,O\}}.$ (23)
We show in Table˜4 the normalized rank across all 4 base models: MxD’s learned experts exhibit no rank deficiencies, providing further evidence of the large potential capacity of MxD layers despite their sparsity constraints on the expert-level.
Table 4: Mean normalized expert matrix rank of Equation˜23 across models for the first 2k experts in $K=32$ trained MxDs – the learned expert matrices are very close to full column rank.
GPT2-124M Pythia-410M Pythia-1.4B Llama-3.2-3B
$9\pm 0.005 0.99±0$$9\pm 0.007$ $9\pm 0.005 0.99±0$$9\pm 0.002$
B.6 Computational cost of sparse layers
We show here that there is minimal difference between the computational cost of sparse layer variants. We first report theoretical layer FLOPs, and then report empirical benchmarks.
Parameter count & FLOPs
We first tabulate in Table˜5 the theoretical parameter counts and inference-time FLOPs for MxDs vs Sparse MLPs. To be consistent with the popular PyTorch library [102] we count one fused multiply-add as one FLOP, and count the Hadamard product between two $d$-dimensional vectors as requiring $d/2$ FLOPs. Note that, For a chosen expert count $N$, we set the width of Sparse MLPs (e.g. Transcoders) to $H:=N+H^{*}$ to parameter-match the models.
Table 5: Here $I,O$ denotes the input and output dimensions, $H^{*}$ the original model’s hidden width, NN the MxD expert count, and $H$ the width of the sparse MLPs.
Parameter count FLOPs
Sparse MLP H(I+O)H\,(I+O) H(I+O)H\,(I+O)
MxD $N\,(I+O)+H^{*}\,(I+O)$ $N\,(I+O)+H^{*}\,(I+O)+O/2$
Empirical benchmarks
We next run benchmarks for a Sparse MLP layer vs MxD with a batch size of 512, and dimensions: $I=H^{*}=O=1024$, and number of experts/features as N=8192,H=9216N=8192,\,H=9216. The results are tabulated in Table˜6, where we see performance (and cost) is similar for the two layers.
Table 6: Here $I,O$ denotes the input and output dimensions, $H^{*}$ the original model’s hidden width, NN the MxD expert count, and $H$ the width of the sparse MLPs.
Peak memory usage (MiB) Latency (ms) Parameter count Reported FLOPs ([102])
Sparse MLP 386.50 1.394 18,874,368 18,874,368
MxD 389.50 1.457 18,874,368 18,874,880
B.7 Sparse probing
Sample-level probing
Here, we follow the SAEBench [19] evaluation protocol. In this ‘sample-level’ setting, each text string is labeled with a binary concept at a global level (e.g., the language of the snippet, or its sentiment). This is in contrast to what we refer to as ‘token-level probing’, where each token within the text samples is labeled individually (e.g., whether a word is a certain part of speech). We perform experiments on a total of $24$ sample-level sparse probing tasks with the same ‘maximum mean difference’ feature filtering applied in [19]. The details of the datasets used are summarized in Table˜7.
Token-level probing
We also explore sparse probing for $10$ features defined at the token-level. For this, we follow [49], and include experiments training probes on the mean feature activations under tokens spanning the surnames of the individuals. We note that this is a significantly harder task, and makes even stronger assumptions about the features the dataset includes, but is nonetheless some additional weak evidence about the relative feature-learning abilities of the sparse models. Through various surnames, we probe for $6$ occupations of individuals, whether or not individuals are alive, and individuals’ labeled gender. We also experimented with probing for compound words as in [49], but found no predictive features in our trained models. Details of the surname token-level probing datasets (and the total training examples the tokenizers could parse) are included in Table˜8.
Table 7: Details of sample-level sparse probing datasets used.
Dataset # Training examples # Test examples Classification task description Number of classes
fancyzhx/ag_news [48] 16,000 4,000 News article topic 4
codeparrot/github-code [103] 20,000 5,000 Programming language 5
amazon_reviews_mcauley_1and5_sentiment [104] 8,000 2,000 Positive/negative review sentiment 2
Helsinki-NLP/europarl [105] 20,000 5,000 European language 5
LabHC/bias_in_bios [106] 32,000 8,000 Profession from bio 8
Table 8: Details of token-level sparse probing datasets used.
Dataset # Training examples # Test examples Classification task description Number of classes
Occupation [49] 4784 1195 Occupation of individual 6
Is alive? [49] 4800 1199 Are they alive 2
Gender [49] 4800 1200 Labeled gender 2
Experimental setup
For sample-level probing, we truncate the input strings to the first $128$ tokens for all datasets but for the Github dataset, where we take the last $128$ tokens to avoid license headers [19, 49]. For token-level probing, we instead take only the last $128$ tokens, where the final token contains the surname of the individual in question in the datasets of [49].
Binary probes are trained on $80\%$ of the training data (randomly shuffled) with the sklearn library’s LogisticRegression module with parameters:
  • class_weight=‘balanced’
  • penalty=‘l2’
  • solver=‘newton-cholesky’
  • max_iter=200
A random seed of $42$ is used throughout the code to ensure reproducibility.
B.7.1 Sparse probing results
We show in Figure˜12 results on $20$ additional (sample-level) sparse probing tasks, where MxDs remain competitive with the baselines. We also plot the expert activation (of the single expert with the highest F1 test set score) for the positive/negative classes for all tasks split across Figures˜13 and 14. One can observe a certain degree of separability between the two semantic clusters of data given by the expert coefficient, thus confirming that individual experts are learning to specialize to particular high-level features.
We also include results on $10$ token-level probing tasks in Figure˜15, with the corresponding activation densities displayed in Figure˜16. Whilst MxDs appear to perform slightly less well here on average, they remain competitive as expected.
(a) Europarl dataset, on GPT2
(b) Europarl dataset, on Pythia-410m
(c) Github code dataset, on GPT2
(d) Github code dataset, on Pythia-410m
(e) Amazon review sentiment dataset, on GPT2
(f) Amazon review sentiment dataset, on Pythia-410m
(g) Bias in Bios dataset, on GPT2
(h) Bias in Bios dataset, on Pythia-410m
Figure 12: Sample-level sparse probing results on individual experts/features; the best F1 score on a held out set is presented.
(a) AG news dataset, on GPT2
(b) AG news dataset, on Pythia-410m
(c) Europarl dataset, on GPT2
(d) Europarl dataset, on Pythia-410m
(e) Github code dataset, on GPT2
(f) Github code dataset, on Pythia-410m
(g) Amazon review sentiment dataset, on GPT2
(h) Amazon review sentiment dataset, on Pythia-410m
Figure 13: [1/2] Sample-level sparse probing results on individual experts for MxDs; here we plot the values of the expert pre-activation for positive/other classes (in the 1-vs-all setting).
(a) Profession from biography, on GPT2
(b) Profession from biography, on Pythia-410m
Figure 14: [2/2] Sample-level sparse probing results on individual experts for MxDs; here we plot the values of the expert pre-activation for positive/other classes (in the 1-vs-all setting).
(a) Occupation surname probing, on GPT2
(b) Occupation surname probing, on Pythia-410m
(c) “Is alive?” surname probing, on GPT2
(d) “Is alive?” surname probing, on Pythia-410m
(e) Gender surname probing, on GPT2
(f) Gender surname probing, on Pythia-410m
Figure 15: Token-level sparse probing results on individual experts/features; the best F1 score on a held out set is presented.
(a) Occupation surname probing, on GPT2
(b) Occupation surname probing, on Pythia-410m
(c) Alive/dead surname probing, on GPT2
(d) Alive/dead surname probing, on Pythia-410m
(e) Gender surname probing, on GPT2
(f) Gender surname probing, on Pythia-410m
Figure 16: Token-level sparse probing results on individual experts for MxDs; here we plot the values of the expert pre-activation for positive/other classes (in the 1-vs-all setting).
B.8 Ablations
We turn next to ablation studies to explore the value of the various model components below:
Figure 17: ReLU+TopK activation function [23] vs ReLU w/ L1 sparsity penalty [22]: both MxDs and Transcoders better recover the cross entropy loss with the TopK activation.
B.8.1 Choice of sparsity constraint
We first train a variety of MxDs on GPT2 models with the TopK activation function [23] and instead train models with a ReLU followed by an explicit $\lambda||.||_{1}$ sparsity penalty on the specialized components in addition to the reconstruction loss [22]. We show the results in Figure˜17, where, similarly to [26], we find the TopK activation to dominate on the sparsity-accuracy frontier–we thus use the TopK activation for all experiments.
Figure 18: Encoder architecture ablation: MSE loss when using ReLU activation vs the GELU used by the base models; and MLPs vs GLUs for Llama (rightmost subfigure).
B.8.2 Choice of MxD encoder
Secondly, we show in Figure˜18 the benefits of MxDs’ flexibility in inheriting the original MLP layer’s encoder form/activation function. All models here are trained from scratch for the same number of tokens and with the same experimental setup as in Section˜3.1, with $K=32$. In the first 3 left-most subfigures, we see the Normalized MSE is as low as half when using GELU vs the non-native ReLU activation.
We next ablate the impact of inheriting the same encoder as the Llama-3.2-3B base model. In the rightmost subfigure of Figure˜18, we train MxDs with ReLU-MLP, GELU-MLP, and Swish-GLU encoders. As can be seen, using a GLU with a Swish activation model (matching the base model architecture) yields a Normalized MSE almost an order of magnitude smaller than MLPs with GELU/ReLU.
Appendix C Feature balance and shared experts
C.1 Expert/feature balance
Following the code of [27, 107], we log how often each unit of specialism/feature is used, over a fixed window of $\sim 1$M tokens. We show in Figure˜19 the feature frequency at the end of training, where we observe that MxDs see a similar healthy balance of experts to the frequency of usage of features in the baselines.
Interestingly, we observe a small peak of experts that fire more frequently in MxDs (e.g., around -2 on the x-axis)–perhaps specializing in common patterns and primitives in natural language.
Figure 19: $\log_{10}$ feature sparsity (following [27, 107]); MxDs’ experts are well-balanced, similar to the baselines’ features.
C.2 Shared experts
We find that, by default, our MxD models naturally learn to use a shared expert, with the remaining experts exhibiting strong specialization in a wide range of themes and linguistic patterns. The use of a shared expert is becoming an increasingly popular design choice, including in the latest Llama 4 models [44]–we therefore allow this pattern to emerge naturally in our base models, further justified through the evidence in [43] that shared experts can enhance specialization among the remaining experts [43]. We highlight, however, that a simple trick of sampling $\hat{K}\sim\text{Unif}\{K-K/a,K+K/a\}$ for the Top-$\hat{K}$ activation at train-time (for e.g. $a:=2$) is sufficient to remove the dominating shared-expert at minimal hit to reconstruction performance, if desired.
We train two sets of models with a base $K=32$ on GPT2-small and pythia-410m, using $a:=2$. We first show in Figure˜20 the indices of the top-activating experts for the 2 model variants on a template prompt, after training has finished. On the left-hand side of Figure˜20, the models route all tokens through the same shared expert at position 1. However, we see on the right-hand side that training with the ‘random-K’ strategy breaks the dependence on a shared expert in position 1. Furthermore, we include in Figure˜21 the corresponding train-time MSE loss for the 4 models here as ablations–observing that the random-K strategy also brings comparable performance. Based on these experiments, we recommend this simple training strategy if one desires MxD models without shared experts.
Expert diversity/collapse
To what extent is the shared expert simply learning the base decoder itself? In an initial attempt to study this, we take the same MxD trained with $K=64$ on GPT2, and tabulate the Frobenius norm of differences between the original model’s decoder $\mathbf{D}^{*}$, our learned base decoder $\mathbf{D}$, and the “shared expert” weight matrix $\mathbf{W}_{\text{shared}}=\mathbf{D}\,\text{diag}(\mathbf{c}_{\text{shared}})$ (with index $\text{shared}=19772$), all of which are matrices of shape $(H\times O)$. As can be seen in Table˜9, their difference is significant, showing that the shared expert does not simply learn the original decoder.
As discussed in the limitations section, we highlight that MxDs as formulated contain no explicit loss terms to discourage expert collapse, or encourage expert balance or diversity. Whilst we find no need for these in our experiments, they may prove useful (or even necessary) in other settings or future work.
Table 9: Distance between learned shared expert, base decoder, and original decoder
$\left\lVert\mathbf{W}_{\text{shared}}-\mathbf{D}^{*}\right\rVert_{F}$ $\left\lVert\mathbf{D}-\mathbf{D}^{*}\right\rVert_{F}$
$\left(\left|\mathbf{W}_{\text{shared}}-\mathbf{D}^{*}\right|\right)_{:3,:3}$
(9 elements of the difference matrix)
208.08 221.6934 $\begin{bmatrix}0.09&0.03&0.06\\ 0.21&0.12&0.07\\ 0.17&0.15&0.11\end{bmatrix}$
Figure 20: Top-activating experts for template prompt with and without using a randomized value of $K$ at train-time for TopK expert selection: randomization largely prevents a shared expert. Shown are the leading 4 tokens and expert indices.
Figure 21: MxD performance with random K sampling: Normalized MSE loss as a function of training steps using a fixed Top $K:=32$ expert selection and when sampling $\hat{K}\sim\text{Unif}\left\{K-\frac{K}{2},K+\frac{K}{2}\right\}$.
Appendix D Detailed experimental setup
We list in Table˜10 the resources used for each experiment: the GPU and the indicative run-time for a single model. The mlp_expansion_factor column refers to the expansion factor applied to the input dimension to generate the MLP width in the sparse layers (i.e. $H:=I\cdot\text{mlp\_expansion\_factor}$).
Table 10: Total training time and resources used to produce the $k=32$ experiments (the required compute being roughly the same across models trained with different $k$).
Model GPU used VRAM Training time d_in mlp_expansion_factor Asset link
GPT2-124m x1 GeForce RTX 3090 24GB 8h 34m 37s 768 32 https://huggingface.co/docs/transformers/en/model_doc/gpt2
Pythia-410m x1 GeForce RTX 3090 24GB 8h 35m 17s 1024 32 https://huggingface.co/EleutherAI/pythia-410m
Pythia-1.4B x1 A100 80GB 23h 25m 23s 2048 32 https://huggingface.co/EleutherAI/pythia-1.4b
Llama-3.2-3B x1 A100 80GB 2d 3m 50s 3072 32 https://huggingface.co/meta-llama/Llama-3.2-3B
D.1 Feature steering details
For the steering experiments, we use two LLM judges to grade generations on two axes. The full template prompt we feed to gemini-2.0-flash and llama-4-scout-17b-16e-instruct is as follows (note that line breaks and emphases are included here only to aid visualization):
Prompt given to LLM judges You are an expert evaluator of synthetic text.
TASK: Rate a collection of {num_samples} samples along two independent axes.
AXIS 1 – CONCEPT COHERENCE:
0.00 no shared concepts/themes/style.
0.25 faint overlap.
0.50 some overlap or similar structure.
0.75 mostly the same concepts or structure; a few partial drifts.
1.00 all snippets clearly share the same concepts, themes, style, or structure.
AXIS 2 – GRAMMATICAL FLUENCY:
0.00 incomprehensible.
0.25 dense errors; meaning often obscured.
0.50 frequent errors; meaning still mostly recoverable.
0.75 minor errors that rarely hinder comprehension.
1.00 completely grammatical and natural.
(Do not penalise fluency if a snippet starts or ends abruptly.).
SCORING: Choose any real value in [0, 1] for each axis.
OUTPUT FORMAT: Respond with exactly two numbers formatted ‘0.00, 0.00’ in the order [coherence, fluency] and no other text or symbols.
TEXT TO EVALUATE: {samples}
Appendix E Additional qualitative results
We show in Figures˜23 and 22 tokens activating the first 9 experts as they appear numerically. We sample 6 bins of expert coefficient value to show both tokens that highly activate the experts and those that do so only mildly. As can be seen, both high- and low-level specializations emerge in both GPT and Pythia models.
Whilst we observe specializations to a range of concepts (such as punctuation, MMO games, words in specific contexts), we do not notice any systemic differences between the types of expert specializations that emerge between the two models in MxD layers.
Figure 22: Tokens activating the first 99 numerical experts on MxDs with $K=32$ trained on Pythia-410m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to MMO games, abbreviations, and words in specific contexts.
Figure 23: Tokens activating the first 99 numerical experts on MxDs with $K=32$ trained on GPT2-124m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to punctuation, names, and months.